Hi
The average mass of a group of boys was 52kg at first. One more boy weighing 84kg joined the group and the average mass became 56kg. Find the number of boys in the group now.
Thanks

Answer by ikleyn(52775)

.

The solution with no words is presented below.


    52*n + 84 = 56*(n+1)


    52n  + 84 = 56n + 56


    84   - 56 = 56n - 52n


       28     = 4n


        n     = 28/4 = 7.    


ANSWER.  There are 7+1 = 8 boys in the group now.

Solved.